True or false: If PQ≅QR , with PQ = 3x – 8, QR = 2x, and RS = 1.5x + 4, then PS = 24.



1
Expert's answer
2021-09-30T17:25:01-0400

Given that

"PQ=3x-8,QR=2x\\RS=1.5x+4,PS=24"

Now

"PQ+QR+RS=PS\\=(3x-8)+2x+(1.5x+4)=24\\=6.5x-4=24\\6.5x=28\\x=4.3\\now\\PQ=3×4.3-8=4.9\\and\\QF=2×4.3=8.6"

Since"PQ{=}\mathllap{/\,}QR"

So the given statement is false